30 June 2013

OSA: David Gonzalez Huerta (LBTO) ISp: John Morris (LBTO) Support: David Thompson, Olga Kuhn, Michelle Edwards (Tucson)

Clouds and rain showers up there - the monsoon is threatening to come, so the telescope has not been open yet...

John Hill and Taras did not have a chance to do their tests on AGw1.

David running some MODS calibrations:
  • bias 8K
  • grpixflats
  • grlamps
  • slitflats through the 0.8", 1" and 1.2" slits.
    • red channel hung...
      • refresh on GUI
      • red reset (and blue reset)
      • aborted script with cntl-c
      • OK thereafter

1 July 2013

OSA: David Gonzalez Huerta (LBTO) ISp: John Morris (LBTO) Support: David Thompson, Olga Kuhn, Michelle Edwards (Tucson)

Clouds, rain. Did not open.

-- OlgaKuhn - 30 Jun 2013
Topic revision: r2 - 17 Jul 2013, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback